# Leray-Schauder Mappings for Operator Learning

### v2

@@ 2. Theoretical Preliminaries
We first prove that
(9) $\inf_{x\in\mathbb{K}}\sum_{j=1}^{n}\mu_{j}(x)>0.$
The same argument is found also in the proof of Theorem 2.2, and we report it here in the $L^{p}$ case to render the two proofs readable independently. By way of contradiction, assume otherwise. Then, let $\{y_{k}\}$ be a sequence in $\mathbb{K}$ such that as $k\longrightarrow+\infty$ we have $\sum_{j=1}^{n}\mu_{j}(y_{k})\longrightarrow 0$. Since $\mathbb{K}$ is a compact in a metric space, it is sequentially compact. Let $y_{k_{r}}$ be a convergent subsequence, and let $y_{k_{r}}\longrightarrow y^{*}$. The set $\{x_{1},\ldots,x_{n}\}$ is an $\epsilon$-net by assumption, therefore we can find $x_{i}$ such that $q:=\|y^{*}-x_{i}\|_{p}<\epsilon$, hence $\epsilon-q>0$. Let $0<\delta<\epsilon-q$. By continuity of the map $\sum_{j=1}^{n}\mu_{i}:\mathbb{K}\longrightarrow\mathbb{R}^{+}_{0}$ we can find a $\delta>0$ such that for all $x\in B(y^{*},\delta)\bigcap\mathbb{K}$ we have $\sum_{j=1}^{n}\mu_{j}(x)\geq\epsilon-q-\delta>0$. Therefore, for $r$ large enough, we have $\sum_{j=1}^{n}\mu_{j}(y_{k_{r}})\geq\epsilon-q-\delta>0$ against the fact that $\sum_{j=1}^{n}\mu_{j}(y_{k})\longrightarrow 0$. This shows that (9) holds true. We put $m:=\inf_{x\in\mathbb{K}}\sum_{j=1}^{n}\mu_{j}(x)$. We also set $M:=\sup_{x\in\mathbb{K}}\sum_{j=1}^{n}\mu_{j}(x)$, which is clearly a finite number since each $\mu_{i}$ is continuous over a compact.
[…]
(10) $\hat{\mu}_{i}(u)=\begin{cases}\epsilon-\mathfrak{I}(|x-x_{i}|^{p})^{\frac{1}{p}},\hskip 28.45274pt\mathfrak{I}(|x-x_{i}|^{p})^{\frac{1}{p}}\leq\epsilon\\ 0,\hskip 102.43008pt\mathfrak{I}(|x-x_{i}|^{p})^{\frac{1}{p}}>\epsilon\end{cases}$
@@
$\leq$ $\sum_{i=1}^{n}\frac{\|x_{i}\|_{p}}{(m-e_{k})m}\cdot\sum_{j=1}^{n}Me_{k}$
$=$ $\frac{n^{2}Me_{k}}{m^{2}-me_{k}}\cdot\max_{i}\|x_{i}\|_{p}.$
@@
Since the functions $\hat{\mu}_{i}:\mathbb{R}^{k}\longrightarrow\mathbb{R}^{+}_{0}$ are continuous with respect to the inputs of the discretization of the integration scheme $\mathfrak{I}$, we can approximate each function $\hat{\mu}_{i}$ by a neural network $F^{i}_{\theta}$, according to the results in [Horn, Fun, LLPS, Lu, Pink] with arbitrarily high precision for each $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$. By thaking the nonlinearity to be a sigmoid, and adding a cutoff at zero we can also ensure that the output of each $F^{i}_{\theta}$ is nonnegative. In particular, we can find architectures such that
@@
for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$. Then, if $k$ is large enough such that $\frac{n^{2}Me_{k}}{m^{2}-me_{k}}\cdot\max_{i}\|x_{i}\|_{p}<\frac{\eta}{2}$, we find that for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$ we have
@@
$\leq$ $\frac{n^{2}Me_{k}}{m^{2}-me_{k}}\cdot\max_{i}\|x_{i}\|_{p}+\sum_{i=1}^{n}\|\frac{\hat{\mu}_{i}(x)x_{i}}{\sum_{j=1}^{n}\hat{\mu}_{j}(x)}-\frac{F^{i}_{\theta}(x)x_{i}}{\sum_{j=1}^{n}F^{j}_{\theta}(x)}\|_{p}$
@@
This completes the proof, once we stack the neural networks $F^{i}_{\theta}$ to obtain $F=(F^{1}_{\theta},\ldots,F^{n}_{\theta})^{T}$. ∎
@@
For $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$ we have that $\|x\|_{\infty}\leq\rho$. Therefore, for a discretization $\{t_{1},\ldots,t_{k}\}$ we have that $\max_{\ell,i}|x(t_{\ell})-x_{i}(t_{\ell})|\leq 2\rho$. It follows that the inputs of $\mathfrak{I}$ are in the bounded hypercube $[-q,q]^{\times k}$, where $q=2^{p}\rho^{p}$. We can therefore apply the results for universal approximation over compacts. ∎
@@
Let $T:X\longrightarrow X$ be continuous map, where $X=L^{p}_{\mu}([a,b])$. Let $\mathbb{K}\subset X$ be a compact such that $T(\mathbb{K})\subset\mathbb{K}$. Then, for any given $\epsilon>0$ we can find $n\in\mathbb{N}$, $\rho>0$, and neural networks $f_{n}$ and $F$ such that
(11) $\|T(x)-\phi_{n}^{-1}f_{n}\phi_{n}(\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)})\|_{p}$ $<$ $\epsilon,$
for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$, where $\bar{B}_{\rho}$ is a radius $\rho$ closed ball in $C^{\kappa,\alpha}([a,b])$, and $g_{i}$ are neural networks for $i=1,\ldots,n$.
@@
The proof is similar to that of Theorem 2.2, where the details following Theorem 2.1 do not depend on the use of the uniform norm. The main difference here lies in the use of Lemma 2.3, to construct a neural network Leray-Schauder mapping $F$ which plays the role of $P_{n}$ in Theorem 2.1, and $\tilde{P}_{n}$ in Theorem 2.2. We therefore provide only the details pertaining to this aspect of the proof. Since $T(\mathbb{K})\subset\mathbb{K}$, we can construct a single Leray-Schauder mapping, as opposed to two as in the proof of Theorem 2.1, where one is used for $\mathbb{K}$, and another one for $T(\mathbb{K})$. The double indices referring to the two mappings can therefore be suppressed.
Observe that by density of smooth functions in $X$, we can assume that an $\epsilon$-net is constructed by taking $\{x_{1},\ldots,x_{n}\}$ smooth functions. Furthermore, since we can approximate each $x_{i}$ by neural networks $g_{i}$ in the uniform norm, we can assume that the $\epsilon$-net consists of neural networks. In order to have $g_{i}$ smooth for each $i$, we can restrict ourselves to smooth (nonpolynomial) nonlinearities.
Then, applying Lemma 2.3 we can find $F:\mathbb{R}^{k}\longrightarrow[0,+\infty)$ such that the Leray-Schauder mapping $P_{n}$ can be approximated by the mapping $x\mapsto\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}$ for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$ with arbitrary precision, where $\rho\geq\max_{i}\|g_{i}\|_{\infty}$. Now, to apply the arguments of Theorem 2.1 and Theorem 2.2 there is one last subtlety to handle. While $P_{n}(\mathbb{K})$ is a compact in $E_{n}$, it is not a priori clear whether the image of $x\mapsto\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}$, for $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$ is compact. However, by construction we have that
@@
$\leq$ $\sum_{i=1}^{n}|\frac{\epsilon+\eta}{m-\eta}|\cdot\|g_{i}\|_{p}$
$\leq$ $\frac{\epsilon}{m-\eta}\max_{i}\|g_{i}\|_{p},$
where we have denoted by $\eta$ the error in the approximation $P_{n}$ via the mapping $x\mapsto\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}$, used the fact that $\mu_{i}(x)\leq\epsilon$ always, and the properties obtained in the proof of Lemma 2.3. Then, the image of the mapping $x\mapsto\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}$ is bounded in the finite dimensional space $E_{n}$, and we can therefore bound it by a closed ball in $E_{n}$ which is compact, since $E_{n}$ with the norm induced by the $p$ norm of $X$ is equivalent to the $n$-dimensional Euclidean space. So, the neural network $f_{n}$ can be defined over a compact, where it is uniformly continuous. Therefore, we can choose the $\eta$ approximation of $P_{n}$ by $x\mapsto\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}$ such that $\|f_{n}(\phi_{n}P_{n}(x))-f_{n}\phi_{n}\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}\|_{p}<\frac{\epsilon}{\|\phi_{n}^{-1}\|}$. Then, we have
$\|T(x)-(\phi_{n}^{-1}f_{n}\phi_{n})\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}\|_{p}$$\leq$ $\|T(x)-TP_{n}(x)\|_{p}+\|TP_{n}(x)-P_{n}TP_{n}(x)\|_{p}$
$+\|P_{n}TP_{n}(x)-(\phi_{n}^{-1}f_{n}\phi_{n})P_{n}(x)\|_{p}$
$+\|(\phi_{n}^{-1}f_{n}\phi_{n})P_{n}(x)-(\phi_{n}^{-1}f_{n}\phi_{n})\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}\|_{p}$
@@
To consider the multivariate case, one needs to use the Hölder space of functions with multiple variables, and needs to substitute cubatures to quadrature rules. Sobolev cubature formulas [Sobolev] can then be used to control the errors in the approximations of the integrations, and as the number of points grow, one can approximate the multidimensional integrals with arbitrary precision.

### v3

@@ 2. Theoretical Construction
As in the case of Theorem 2.2, letting $\mu_{i}$ indicate the maps used to define $P_{n}$, we have
(9) $\min_{x\in\mathbb{K}}\sum_{j=1}^{n}\mu_{j}(x)>0.$
We put $m:=\min_{x\in\mathbb{K}}\sum_{j=1}^{n}\mu_{j}(x)$. We also set $M:=\max_{x\in\mathbb{K}}\sum_{j=1}^{n}\mu_{j}(x)$, which is a finite number since each $\mu_{i}$ is continuous over a compact.
We now introduce an integration scheme such that we can approximate the integrals that give $\|\cdot\|_{p}$ by a quadrature rule uniformly on elements of $\mathbb{K}$ that also lie in a given Hölder ball centered at the origin, $\bar{B}_{\rho}$, as in the statement of the lemma. For simplicity, we can think of this as being the forward rectangle rule, even though the same argument can be applied to other rules that converge faster (e.g. Trapezoidal, Cavalieri-Simpson, Boole) upon changing the error bounds, and taking $k$ large enough in the statement of this lemma, depending on the value of $\kappa$. A similar procedure was used in [uni] to approximate integral operators with applications to the transformer architecture. For $u\in\bar{B}_{\rho}$, we have that $\sup_{[a,b]}|u^{\prime}|\leq\rho$, and therefore we find that the quadrature rule gives uniform error bounds of $e_{k}=\frac{b-a}{2k}\rho$ for all $u\in\mathbb{K}\bigcap\bar{B}_{\rho}$, where $k$ is the number of points used to discretize the interval $[a,b]$. We indicate the quadrature rule to approximate the integral $\int_{a}^{b}fd\mu$ by $\mathfrak{I}(f)$. In addition, as $k\longrightarrow+\infty$ we have that $|\hat{\mu}_{i}(x)-\mu_{i}(x)|\longrightarrow 0$ for all $i=1,\ldots,n$ and for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$, where we define
(10) $\hat{\mu}_{i}(x)=\begin{cases}\epsilon-\mathfrak{I}(|x-x_{i}|^{p})^{\frac{1}{p}},\hskip 28.45274pt\mathfrak{I}(|x-x_{i}|^{p})^{\frac{1}{p}}\leq\epsilon\\ 0,\hskip 102.43008pt\mathfrak{I}(|x-x_{i}|^{p})^{\frac{1}{p}}>\epsilon\end{cases}$
@@
$\leq$ $\sum_{i=1}^{n}\frac{\|x_{i}\|_{p}}{(m-e_{k})m}\cdot\sum_{j=1}^{n}2Me_{k}$
$=$ $\frac{2n^{2}Me_{k}}{m^{2}-me_{k}}\cdot\max_{i}\|x_{i}\|_{p}.$
@@
Since the functions $\hat{\mu}_{i}:\mathbb{R}^{k}\longrightarrow\mathbb{R}^{+}_{0}$ are continuous with respect to the inputs of the discretization of the integration scheme $\mathfrak{I}$, we can approximate each function $\hat{\mu}_{i}$ by a neural network $F^{i}_{\theta}$, according to the results in [Horn, Fun, LLPS, Lu, Pink] with arbitrarily high precision for each $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$. In particular, we can find architectures such that
@@
for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$. Then, if $k$ is large enough such that $\frac{2n^{2}Me_{k}}{m^{2}-me_{k}}\cdot\max_{i}\|x_{i}\|_{p}<\frac{\eta}{2}$, we find that for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$ we have
@@
$\leq$ $\frac{2n^{2}Me_{k}}{m^{2}-me_{k}}\cdot\max_{i}\|x_{i}\|_{p}+\sum_{i=1}^{n}\|\frac{\hat{\mu}_{i}(x)x_{i}}{\sum_{j=1}^{n}\hat{\mu}_{j}(x)}-\frac{F^{i}_{\theta}(x)x_{i}}{\sum_{j=1}^{n}F^{j}_{\theta}(x)}\|_{p}$
@@
This completes the proof. ∎
Remark 2.5.
We notice that upon taking some suitable nonlinearity, we can also ensure that the output of each $F^{i}_{\theta}$ in the proof of Lemma 2.4 is nonnegative. In this sense, the map $x\mapsto(\frac{F^{1}_{\theta}(x)}{\sum_{j=1}^{n}F^{j}_{\theta}(x)},\ldots,\frac{F^{n}_{\theta}(x)}{\sum_{j=1}^{n}F^{j}_{\theta}(x)})^{T}$ resembles the components of a Leray-Schauder map $P_{n}$.
@@
For $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$ we have that $\|x\|_{\infty}\leq\rho$. Therefore, for a discretization $\{t_{1},\ldots,t_{k}\}$ we have that $\max_{\ell,i}|x(t_{\ell})-x_{i}(t_{\ell})|\leq 2\rho$. It follows that the inputs of $\mathfrak{I}$ are in the bounded hypercube $q,q]^{\times k}$, where $q=2^{p}\rho^{p}$. We can therefore apply the results for universal approximation over compacts. ∎
In the following result, we consider nonlinearities for the neural networks that are sufficiently smooth, in the sense defined above.
@@
Let $T:X\longrightarrow X$ be a continuous map, where $X=L^{p}_{\mu}([a,b])$. Let $\mathbb{K}\subset X$ be a compact in $X$. Then, for any given $\epsilon>0$ we can find $n,m\in\mathbb{N}$, $\rho>0$, and neural networks $\{g_{i}\}_{i=1}^{n},\{h_{j}\}_{j=1}^{m},f_{n,m}$ and $\{F^{i}\}_{i=1}^{n}$ such that
(11) $\|T(x)-\phi_{m}^{-1}f_{n,m}\phi_{n}(\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)})\|_{p}$ $<$ $\epsilon,$
for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$, where $\bar{B}_{\rho}$ is a radius $\rho$ closed ball in $C^{\kappa+1,\alpha}([a,b])$, and the neural networks $h_{i}$ span the domain of $\phi_{m}$.
@@
The proof is similar to that of Theorem 2.2, where the details following Theorem 2.1 do not depend on the use of the uniform norm. The main difference here lies in the use of Lemma 2.4, to construct a neural network Leray-Schauder mapping $F$ which plays the role of $P_{n}$ in Theorem 2.1, and $\tilde{P}_{n}$ in Theorem 2.2. We therefore provide only the details pertaining to this aspect of the proof. In the construction of $P_{n}$, we find continuous functions $\{x_{1},\ldots,x_{n}\}$ such that $\mathbb{K}$ is contained in $\delta$-balls centered at $x_{i}$. Furthermore, since we can approximate each $x_{i}$ by neural networks $g_{i}$ in the uniform norm, we can assume that the each element used to construct $P_{n}$ is a neural network, and upon taking sufficiently smooth nonlinearities, we can assume that they are in $C^{\kappa+1,\alpha}([a,b])$, so that the function $|g_{i}|^{p}$ is in $C^{\kappa,\alpha}([a,b])$, following essentially the same procedure as in Theorem 2.2. We assume that the $g_{i}$’s span an $n$-dimensional space to simplify notation, since the proof is adapted directly as in [EZ_proj] if otherwise. Then, applying Lemma 2.4 we can find neural networks $F^{i}$ such that the Leray-Schauder mapping $P_{n}$ can be approximated by the mapping $x\mapsto\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}$ for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$ with arbitrary precision, where $\rho\geq\max_{i}\|g_{i}\|_{\infty}$. Now, to apply the arguments of Theorem 2.1 and Theorem 2.2 there is one last subtlety to handle. While $P_{n}(\mathbb{K})$ is a compact in $E_{n}$, it is not a priori clear whether the image of $x\mapsto\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}$, for $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$ is compact. However, by construction we have that
@@
$\leq$ $\sum_{i=1}^{n}|\frac{\delta+\eta}{m_{0}-\eta}|\cdot\|g_{i}\|_{p}$
$\leq$ $n\frac{\delta+\eta}{m_{0}-\eta}\max_{i}\|g_{i}\|_{p},$
where we have denoted by $\eta$ the error in the approximation $P_{n}$ via the mapping $x\mapsto\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}$, we used the fact that $\mu_{i}(x)\leq\delta$ always, we have denoted by $m_{0}$ the minimum of $P_{n}$ on $\mathbb{K}$, and we used the properties obtained in the proof of Lemma 2.4. Then, the image of the mapping $x\mapsto\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}$ is bounded in the finite dimensional space $E_{n}$ for all $\eta$ sufficiently small. We can therefore bound it by a closed ball in $E_{n}$ which is compact, since $E_{n}$ with the norm induced by the $p$ norm of $X$ is equivalent to the $n$-dimensional Euclidean space. So, the neural network $f_{n,m}$ takes inputs for the values of $P_{n}(\mathbb{K})$ and the image of $x\mapsto\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}$ over a compact, where it is uniformly continuous. Therefore, we can choose the $\eta$ approximation of $P_{n}$ by $x\mapsto\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}$ such that $\|f_{n,m}(\phi_{n}P_{n}(x))-f_{n,m}\phi_{n}\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}\|_{p}<\frac{\epsilon}{\|\phi_{m}^{-1}\|}$. Then, we have
$\|T(x)-(\phi_{m}^{-1}f_{n}\phi_{n})\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}\|_{p}$ $\leq$ $\|T(x)-TP_{n}(x)\|_{p}+\|TP_{n}(x)-P_{m}TP_{n}(x)\|_{p}$
$+\|P_{m}TP_{n}(x)-(\phi_{m}^{-1}f_{n,m}\phi_{n})P_{n}(x)\|_{p}$
$+\|(\phi_{m}^{-1}f_{n,m}\phi_{n})P_{n}(x)-(\phi_{m}^{-1}f_{n,m}\phi_{n})\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)}\|_{p}$
@@
We notice that the proof can proceed in the same way if $T:X\longrightarrow Y$ where $Y=L^{q}_{\mu}([a,b])$ with $q$ possibly different from $p$.
We now discuss the generalization of the previous result to the multivariate case $L^{p}([a,b]^{r})$, in which we will replace the quadrature rules in Lemma 2.4 by cubature formulas [Sobolev], restricting ourselves to rules that are continuous in their inputs. We consider the space $C^{\kappa+1,\alpha}([a,b]^{r})$, on $a,b]^{r}$, and assume as before $p$ sufficiently large. As in the previous result, the approximation holds for those functions in $\mathbb{K}$ with some extra regularity.
Theorem 2.9.
Let $T:X\longrightarrow Y$ be continuous map, where $X=L^{p_{1}}_{\mu}([a,b]^{r_{1}})$ and $Y=L^{p_{2}}_{\mu}([c,d]^{r_{2}})$. Let $\mathbb{K}\subset X$ be a compact in $X$. Then, for any given $\epsilon>0$ we can find $n,m\in\mathbb{N}$, $\rho>0$, and neural networks $\{g_{i}\}_{i=1}^{n},\{h_{j}\}_{j=1}^{m},f_{n,m}$ and $\{F^{i}\}_{i=1}^{n}$ such that
(12) $\|T(x)-\phi_{m}^{-1}f_{n,m}\phi_{n}(\sum_{i=1}^{n}\frac{F^{i}(x)g_{i}}{\sum_{j=1}^{n}F^{j}(x)})\|_{p_{2}}$ $<$ $\epsilon,$
for all $x\in\mathbb{K}\bigcap\bar{B}_{\rho}$, where $B_{\rho}$ is a radius $\rho$ ball in $C^{\kappa+1,\alpha}([a,b]^{r})$ with $p_{1}\geq\kappa+\alpha$, and the neural networks $h_{i}$ span the domain of $\phi_{m}$.
Proof.
The proof proceeds in the same way as for Theorem 2.7, with the main difference lying in Lemma 2.4. In this case we use cubature rules to construct maps $\hat{\mu}_{i}$ to approximate the operators $P_{n}$. Restricting to those elements of $\mathbb{K}$ that are also in $\bar{B}_{\rho}$ guarantees that by increasing the grid points used for the cubature formulas, the approximation $\hat{\mu}_{i}$ can be made arbitrarily precise. Moreover, the inputs of the cubatures lie in some hypercube of the Euclidean space, so that the neural networks can be chosen to approximate the cubaature formulas on a compact. ∎